OpenStudy (anonymous):

Circle O has a center at (-2, -2) and a diameter of 10 units. Which point lies on circle O? f (8, 8) g (-6, -5) h (-2, -2) j (6, 4)

OpenStudy (anonymous):

\[\left( x+2 \right)^2+\left( y+2 \right)^2=5^2\] plug each point in this eq. and see which satisfies it.

OpenStudy (anonymous):

thank you! @surjithayer